# Simple Estimation of Semiparametric Models with Measurement Errors

### v3

@@ Appendix D Details on the sufficient condition (17)
where the last equality follows from (17). $\blacksquare$
Appendix E Evaluating the remaining bias of the MERM estimator
In this section, we provide additional details on the derivation of the (higher-order) bias of the MERM estimator of $v^{\prime}\hat{\beta}$ and discuss potential implications for inference.
The derivation of the expression for $\text{Bias}(v^{\prime}\hat{\beta})$ provided in equation (18) is based on the standard expansion of the GMM first order conditions around $\beta_{0}$. Following the lines of the proof of Lemma 1, we can consider a higher order expansion of $\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})$ to the order $o(\tau^{\overline{K}})$, which implies
$\mathbb{E}\left[\psi(X_{i},S_{i},\theta_{0},\gamma_{0n})\right]\approx\gamma_{0\overline{K}}\mathbb{E}\left[g_{x}^{(\overline{K})}(X_{i},S_{i},\theta_{0})\right],$
where $\overline{K}=K+2$ if $K$ is even and $\varepsilon_{i}$ is symmetric, and $\overline{K}=K+1$ otherwise. Hence,
$\text{Bias}(v^{\prime}\hat{\beta})\approx-v^{\prime}(\Psi^{\prime}\Xi\Psi)^{-1}\Psi^{\prime}\Xi\;\mathbb{E}\left[g_{x}^{(\overline{K})}(X_{i},S_{i},\theta_{0})\right]\gamma_{0\overline{K}}.$ (E.1)
Notice that, except for $\gamma_{0\overline{K}}$, all the remaining elements on the right hand side of (E.1) are either known or can be consistently estimated. The unknown $\gamma_{0\overline{K}}$ is given by the recursive relationship (12), and the only unknown determinant of $\gamma_{0\overline{K}}$ is $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ because the lower order moments of $\varepsilon_{i}$ can be recovered from $\hat{\gamma}$ with a required precision. Hence, the ranges of the possible values of $\gamma_{0\overline{K}}$ and, subsequently, of $\text{Bias}(v^{\prime}\hat{\beta})$ can be determined based on an a priori range of possible values of $\mathbb{E}[\varepsilon_{i}^{\overline{K}}]$ (or, equivalently, of $\mathbb{E}[(\varepsilon_{i}/\sigma_{\varepsilon})^{\overline{K}}]$ combined with the estimated $\sigma_{\varepsilon}^{2}$) specified by the researcher.
Example (Evaluating the bias with $K=2$).
Suppose the researcher uses the MERM estimator with $K=2$, and believes that $\mathbb{E}\left[\varepsilon_{i}^{3}\right]=0$. Then, the bound on the bias depends on the $\mathbb{E}\left[\varepsilon_{i}^{4}\right]$. The variance $\sigma_{\varepsilon}^{2}$ is estimated by the MERM estimator, hence it is convenient to bound the kurtosis of $\varepsilon_{i}$ as $\mathbb{E}\left[\varepsilon_{i}^{4}\right]/\sigma_{\varepsilon}^{4}\in[1,\overline{\kappa}]$ for some $\overline{\kappa}$. For example, one could take $\overline{\kappa}=10$ as a conservative bound for the kurtosis. [^13] For Gaussian $\varepsilon_{i}$ the kurtosis $\kappa$ is $3$. Student’s $t(\nu)$ distribution has $\kappa<10$ for all $\nu\geq 5$ (the kurtosis and the $4$th moment do not exist for $\nu\leq 4$). Since $\hat{\gamma}$ estimates $\gamma_{0}=\gamma_{02}=\sigma_{\varepsilon}^{2}/2$, the range of possible values of $\gamma_{04}$ can be approximately bounded using
$\gamma_{04}=\frac{\mathbb{E}\left[\varepsilon_{i}^{4}\right]-6\sigma_{\varepsilon}^{4}}{24}\in\left[-\frac{5}{24}\sigma_{\varepsilon}^{4},\;\frac{\overline{\kappa}-6}{24}\sigma_{\varepsilon}^{4}\right]=\left[-\frac{5}{6}\gamma_{02}^{2},\;\frac{\overline{\kappa}-6}{6}\gamma_{02}^{2}\right].{}$
Hence, the range of possible values of $\text{Bias}(v^{\prime}\hat{\beta})$ can be approximated as
$\text{Bias}(v^{\prime}\hat{\beta})\in\begin{cases}\left[-5\overline{b}\hat{\gamma}^{2}/6,\;\overline{b}\hat{\gamma}^{2}(\overline{\kappa}-6)/6\right],&\text{if $\overline{b}\geq 0$},\\ \left[\overline{b}\hat{\gamma}^{2}(\overline{\kappa}-6)/6,\;-5\overline{b}\hat{\gamma}^{2}/6\right],&\text{otherwise},\end{cases}$
where
$\overline{b}\equiv-v^{\prime}(\hat{\Psi}^{\prime}\hat{\Xi}\hat{\Psi})^{-1}\hat{\Psi}^{\prime}\hat{\Xi}\;\left(\frac{1}{n}\sum_{i=1}^{n}g_{x}^{(4)}(X_{i},S_{i},\hat{\theta})\right).$
Notice that if $\overline{\kappa}\leq 6$, we have $\gamma_{04}\leq 0$, and the sign of $\text{Bias}(v^{\prime}\hat{\beta})$ becomes known.
If (the absolute value of the worst case) $\text{Bias}(v^{\prime}\hat{\beta})$ is sufficiently small relative to its standard error $\sqrt{v^{\prime}\hat{\Sigma}v/n}$, the estimator $v^{\prime}\hat{\beta}$ is approximately unbiased and the inference on $v^{\prime}\beta_{0}$ can be based on equation (14). Otherwise, the researcher could consider using a higher $K$ and/or possibly employing inference tools that take into account the remaining bias (e.g., ArmstrongKolesar2021QE).
Appendix F MERM derivation when $\sigma_{\varepsilon}$ is not small
@@
$\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta_{0})]=0,\quad X_{i}=X_{i}^{\ast}+\varepsilon_{i},\quad\mathbb{E}[\varepsilon_{i}]=0.$ (F.1)
@@
Then, we can rewrite equation (F.1) as
@@
$\widetilde{\psi}(\widetilde{X}_{i},S_{i},\theta,\widetilde{\gamma})=\widetilde{g}(\widetilde{X}_{i},S_{i},\theta)-\sum_{k=2}^{K}\widetilde{\gamma}_{k}\widetilde{g}_{x}^{(k)}(\widetilde{X}_{i},S_{i},\theta),$ (F.2)
@@
Finally, let us now show that moment condition $\widetilde{\psi}$ in equation (F.2) is numerically identical to $\psi$ in equation (11) with $\gamma_{k}=\sigma_{X^{\ast}}^{k}\widetilde{\gamma}_{k}$. Note that for $\widetilde{x}=x/\sigma_{X^{\ast}}$ we have $\widetilde{g}_{\widetilde{x}}^{(k)}(\widetilde{x},s,\theta)\equiv\nabla_{\widetilde{x}}^{k}g(\sigma_{X^{\ast}}\widetilde{x},s,\theta)=\sigma_{X^{\ast}}^{k}g_{a}^{(k)}(a,s,\theta)|_{a=\sigma_{X^{\ast}}\widetilde{x}}=\sigma_{X^{\ast}}^{k}g_{x}^{(k)}(x,s,\theta)$, and hence
$\widetilde{\psi}(\widetilde{X}_{i},S_{i},\theta,\widetilde{\gamma})$ $=$ $g(\sigma_{X^{\ast}}\widetilde{X}_{i},S_{i},\theta)-\sum_{k=2}^{K}\left(\widetilde{\gamma}_{k}\sigma_{X^{\ast}}^{k}\right)g_{x}^{(k)}(\sigma_{X^{\ast}}\widetilde{X}_{i},S_{i},\theta)$
@@
Appendix G Some Implementation Details

### v4

@@ Appendix E Details on the sufficient condition (17)
where the last equality follows from (17).
Appendix F Additional Numerical Results
In this section, we provide additional numerical results for the experiments considered in Sections 3.2 and 4.1.
F.1 Additional numerical results for Section 3.2
In this section, we provide additional simulation results for the same experiment as considered in Section 3.2 but for larger values of $\tau\in\{1,1.5,2\}$. The results are provided in Table 6 below, reporting the same statistics as in Table LABEL:tab:_MC_mult_logit_2 in the main text.
Table 6: Simulation results for the multinomial logit model
MLE $K=2$ $K=4$
bias, $10^{-2}$ std, $10^{-2}$ rmse, $10^{-2}$ size bias, $10^{-2}$ std, $10^{-2}$ rmse, $10^{-2}$ size bias, $10^{-2}$ std, $10^{-2}$ rmse, $10^{-2}$ size
$\tau=1$
$\partial p_{1}/\partial x$ -16.14 0.70 16.15 100.00 -11.45 2.21 11.66 99.16 0.21 3.53 3.54 5.44
$\partial p_{1}/\partial w_{1}$ 11.82 1.40 11.91 100.00 8.49 2.46 8.84 96.14 0.48 2.88 2.92 6.30
$\partial p_{1}/\partial w_{2}$ 2.08 0.67 2.19 89.28 1.63 0.84 1.84 56.78 0.10 1.04 1.05 5.94
$\partial p_{2}/\partial x$ 8.78 0.65 8.80 100.00 6.59 1.50 6.76 97.64 -0.04 2.53 2.53 6.42
$\partial p_{2}/\partial w_{1}$ -5.87 0.71 5.91 100.00 -4.24 1.23 4.42 96.08 -0.24 1.44 1.46 6.26
$\partial p_{2}/\partial w_{2}$ -4.03 1.27 4.22 89.50 -3.20 1.62 3.59 57.34 -0.18 2.09 2.10 6.24
$\partial p_{0}/\partial x$ 7.36 0.61 7.39 100.00 4.86 1.33 5.04 94.72 -0.17 1.95 1.96 5.30
$\partial p_{0}/\partial w_{1}$ -5.95 0.73 6.00 100.00 -4.24 1.24 4.42 96.14 -0.24 1.44 1.46 6.28
$\partial p_{0}/\partial w_{2}$ 1.94 0.61 2.04 89.52 1.57 0.79 1.76 57.96 0.08 1.05 1.05 6.24
$\tau=3/2$
$\partial p_{1}/\partial x$ -19.01 0.50 19.01 100.00 -16.59 1.32 16.64 100.00 -2.29 3.67 4.33 16.56
$\partial p_{1}/\partial w_{1}$ 14.07 1.34 14.13 100.00 12.83 1.90 12.97 99.98 2.70 3.40 4.34 24.14
$\partial p_{1}/\partial w_{2}$ 2.31 0.64 2.40 95.70 2.20 0.75 2.32 87.72 0.56 1.18 1.31 12.18
$\partial p_{2}/\partial x$ 9.97 0.47 9.98 100.00 9.05 0.95 9.10 100.00 1.58 2.70 3.13 16.10
$\partial p_{2}/\partial w_{1}$ -6.96 0.69 6.99 100.00 -6.40 0.96 6.48 99.98 -1.35 1.71 2.18 24.08
$\partial p_{2}/\partial w_{2}$ -4.44 1.21 4.60 95.88 -4.29 1.43 4.52 87.88 -1.09 2.35 2.59 12.68
$\partial p_{0}/\partial x$ 9.04 0.46 9.05 100.00 7.54 0.94 7.60 100.00 0.71 2.05 2.18 9.46
$\partial p_{0}/\partial w_{1}$ -7.11 0.71 7.15 100.00 -6.42 0.98 6.50 99.98 -1.35 1.70 2.17 23.90
$\partial p_{0}/\partial w_{2}$ 2.13 0.58 2.21 95.90 2.09 0.69 2.20 88.08 0.53 1.17 1.29 13.00
$\tau=2$
$\partial p_{1}/\partial x$ -20.28 0.38 20.29 100.00 -18.86 0.88 18.88 100.00 -5.98 3.91 7.15 55.08
$\partial p_{1}/\partial w_{1}$ 15.09 1.32 15.15 100.00 14.91 1.66 15.00 100.00 5.87 3.87 7.03 55.08
$\partial p_{1}/\partial w_{2}$ 2.39 0.63 2.47 97.42 2.37 0.71 2.47 94.38 1.12 1.30 1.72 26.44
$\partial p_{2}/\partial x$ 10.45 0.36 10.46 100.00 9.97 0.69 9.99 100.00 3.80 2.82 4.73 47.94
$\partial p_{2}/\partial w_{1}$ -7.45 0.68 7.48 100.00 -7.43 0.85 7.48 100.00 -2.95 1.95 3.54 54.96
$\partial p_{2}/\partial w_{2}$ -4.58 1.18 4.73 97.42 -4.61 1.36 4.81 94.54 -2.20 2.54 3.36 27.46
$\partial p_{0}/\partial x$ 9.83 0.36 9.84 100.00 8.88 0.70 8.91 100.00 2.18 2.21 3.11 28.50
$\partial p_{0}/\partial w_{1}$ -7.64 0.71 7.68 100.00 -7.48 0.87 7.53 100.00 -2.92 1.93 3.50 54.80
$\partial p_{0}/\partial w_{2}$ 2.19 0.57 2.26 97.42 2.24 0.66 2.34 94.62 1.08 1.26 1.65 27.78
  • This table reports the simulated finite sample bias, standard deviation, RMSE, and size of the MLE and the MERM estimators and the corresponding t-tests for the partial derivatives $\partial p_{j}(x,w,\theta_{0})/\partial x$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{1}$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{2}$ for $j\in\{1,2,0\}$ evaluated at the population mean. The true values of the marginal effects are $(\partial p_{1}/\partial x,\partial p_{2}/\partial x,\partial p_{0}/\partial x)=(0.222,-0.111,-0.111)$ and zeros for the rest. The results are based on 5,000 replications.
F.2 Additional numerical results for Section 4.1
In this section, we provide additional numerical results for the experiment considered in Section 4.1. In particular, Tables 7 and 8 below report the same statistics as Table 5 in the main text, but for smaller and larger sample sizes $n=1000$ and $n=4000$.
Table 7: Choice of $K$ simulation results for the multinomial logit model, $n=1000$
MLE $K=2$ $K=4$ data-driven $K$
bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$
$\tau=1/4$
$\partial p_{1}/\partial x$ -3.22 3.74 1.34 3.96 2.10 4.56 1.37 4.05
$\partial p_{1}/\partial w_{1}$ 2.29 3.27 -0.13 3.32 -0.57 3.39 -0.15 3.34
$\partial p_{1}/\partial w_{2}$ 0.52 1.19 -0.01 1.22 -0.10 1.26 -0.02 1.23
$\partial p_{2}/\partial x$ 1.93 2.55 -0.73 2.78 -1.18 3.11 -0.75 2.82
$\partial p_{2}/\partial w_{1}$ -1.15 1.64 0.07 1.66 0.29 1.70 0.07 1.67
$\partial p_{2}/\partial w_{2}$ -1.02 2.36 0.03 2.46 0.21 2.53 0.04 2.46
$\partial p_{0}/\partial x$ 1.28 1.93 -0.61 2.16 -0.92 2.41 -0.62 2.20
$\partial p_{0}/\partial w_{1}$ -1.14 1.63 0.07 1.66 0.29 1.70 0.07 1.67
$\partial p_{0}/\partial w_{2}$ 0.51 1.17 -0.02 1.24 -0.11 1.27 -0.02 1.24
$\tau=1/2$
$\partial p_{1}/\partial x$ -8.97 9.10 -1.82 4.10 1.76 4.57 1.35 4.82
$\partial p_{1}/\partial w_{1}$ 6.42 6.79 1.88 4.04 -0.35 3.52 -0.10 3.73
$\partial p_{1}/\partial w_{2}$ 1.32 1.67 0.42 1.33 -0.05 1.32 0.00 1.34
$\partial p_{2}/\partial x$ 5.21 5.39 1.16 2.91 -0.97 3.20 -0.73 3.30
$\partial p_{2}/\partial w_{1}$ -3.21 3.39 -0.94 2.02 0.18 1.76 0.05 1.87
$\partial p_{2}/\partial w_{2}$ -2.59 3.26 -0.83 2.65 0.11 2.65 0.01 2.68
$\partial p_{0}/\partial x$ 3.76 3.95 0.66 2.18 -0.79 2.50 -0.62 2.56
$\partial p_{0}/\partial w_{1}$ -3.21 3.40 -0.94 2.02 0.18 1.76 0.05 1.86
$\partial p_{0}/\partial w_{2}$ 1.27 1.60 0.41 1.33 -0.06 1.33 -0.01 1.35
$\tau=3/4$
$\partial p_{1}/\partial x$ -13.36 13.41 -7.17 7.97 1.07 4.74 1.07 4.75
$\partial p_{1}/\partial w_{1}$ 9.68 9.90 5.59 6.67 0.17 3.79 0.17 3.79
$\partial p_{1}/\partial w_{2}$ 1.85 2.09 1.15 1.69 0.07 1.41 0.07 1.41
$\partial p_{2}/\partial x$ 7.48 7.57 4.26 4.91 -0.54 3.36 -0.54 3.36
$\partial p_{2}/\partial w_{1}$ -4.82 4.94 -2.80 3.34 -0.09 1.90 -0.09 1.90
$\partial p_{2}/\partial w_{2}$ -3.59 4.05 -2.27 3.33 -0.13 2.82 -0.13 2.82
$\partial p_{0}/\partial x$ 5.87 5.96 2.91 3.53 -0.53 2.67 -0.53 2.68
$\partial p_{0}/\partial w_{1}$ -4.86 4.97 -2.79 3.34 -0.08 1.90 -0.08 1.90
$\partial p_{0}/\partial w_{2}$ 1.74 1.96 1.12 1.64 0.06 1.41 0.06 1.42
  • This table reports the simulated finite sample bias and RMSE of the MLE and the MERM estimators for the partial derivatives $\partial p_{j}(x,w,\theta_{0})/\partial x$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{1}$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{2}$ for $j\in\{1,2,0\}$ evaluated at the population mean. The true values of the marginal effects are $(\partial p_{1}/\partial x,\partial p_{2}/\partial x,\partial p_{0}/\partial x)=(0.222,-0.111,-0.111)$ and zeros for the rest. The results are based on 5,000 replications.
Table 8: Choice of $K$ simulation results for the multinomial logit model, $n=4000$
MLE $K=2$ $K=4$ data-driven $K$
bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$ bias, $10^{-2}$ rmse, $10^{-2}$
$\tau=1/4$
$\partial p_{1}/\partial x$ -3.28 3.41 0.39 1.90 0.56 1.93 0.39 1.90
$\partial p_{1}/\partial w_{1}$ 2.32 2.57 -0.08 1.59 -0.15 1.57 -0.08 1.59
$\partial p_{1}/\partial w_{2}$ 0.48 0.72 -0.04 0.62 -0.05 0.61 -0.04 0.62
$\partial p_{2}/\partial x$ 1.97 2.13 -0.23 1.35 -0.33 1.37 -0.23 1.35
$\partial p_{2}/\partial w_{1}$ -1.16 1.28 0.04 0.80 0.08 0.78 0.04 0.79
$\partial p_{2}/\partial w_{2}$ -0.96 1.43 0.07 1.23 0.11 1.23 0.07 1.23
$\partial p_{0}/\partial x$ 1.31 1.49 -0.16 1.01 -0.23 1.03 -0.16 1.01
$\partial p_{0}/\partial w_{1}$ -1.16 1.29 0.04 0.80 0.08 0.78 0.04 0.80
$\partial p_{0}/\partial w_{2}$ 0.47 0.71 -0.04 0.62 -0.05 0.62 -0.04 0.62
$\tau=1/2$
$\partial p_{1}/\partial x$ -9.00 9.03 -1.52 2.41 0.49 2.03 0.48 2.04
$\partial p_{1}/\partial w_{1}$ 6.43 6.52 1.06 1.99 -0.11 1.65 -0.10 1.65
$\partial p_{1}/\partial w_{2}$ 1.28 1.38 0.21 0.68 -0.04 0.65 -0.04 0.65
$\partial p_{2}/\partial x$ 5.22 5.27 0.90 1.64 -0.30 1.46 -0.29 1.47
$\partial p_{2}/\partial w_{1}$ -3.21 3.25 -0.53 1.00 0.05 0.82 0.05 0.83
$\partial p_{2}/\partial w_{2}$ -2.51 2.70 -0.42 1.35 0.09 1.30 0.09 1.30
$\partial p_{0}/\partial x$ 3.78 3.83 0.62 1.19 -0.20 1.10 -0.19 1.10
$\partial p_{0}/\partial w_{1}$ -3.22 3.26 -0.53 1.00 0.05 0.82 0.05 0.83
$\partial p_{0}/\partial w_{2}$ 1.23 1.32 0.21 0.67 -0.05 0.65 -0.05 0.65
$\tau=3/4$
$\partial p_{1}/\partial x$ -13.37 13.38 -6.49 6.79 0.41 2.25 0.41 2.25
$\partial p_{1}/\partial w_{1}$ 9.68 9.73 4.44 4.84 -0.02 1.80 -0.02 1.80
$\partial p_{1}/\partial w_{2}$ 1.80 1.87 0.91 1.12 -0.03 0.69 -0.03 0.69
$\partial p_{2}/\partial x$ 7.48 7.50 3.83 4.08 -0.24 1.63 -0.24 1.63
$\partial p_{2}/\partial w_{1}$ -4.82 4.85 -2.22 2.42 0.01 0.90 0.01 0.90
$\partial p_{2}/\partial w_{2}$ -3.50 3.63 -1.80 2.22 0.06 1.39 0.06 1.39
$\partial p_{0}/\partial x$ 5.89 5.91 2.66 2.87 -0.16 1.22 -0.16 1.22
$\partial p_{0}/\partial w_{1}$ -4.86 4.89 -2.22 2.42 0.01 0.90 0.01 0.90
$\partial p_{0}/\partial w_{2}$ 1.70 1.76 0.89 1.09 -0.03 0.69 -0.03 0.69
  • This table reports the simulated finite sample bias and RMSE of the MLE and the MERM estimators for the partial derivatives $\partial p_{j}(x,w,\theta_{0})/\partial x$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{1}$, $\partial p_{j}(x,w,\theta_{0})/\partial w_{2}$ for $j\in\{1,2,0\}$ evaluated at the population mean. The true values of the marginal effects are $(\partial p_{1}/\partial x,\partial p_{2}/\partial x,\partial p_{0}/\partial x)=(0.222,-0.111,-0.111)$ and zeros for the rest. The results are based on 5,000 replications.
Appendix G MERM derivation when $\sigma_{\varepsilon}$ is not small
@@
$\mathbb{E}[g(X_{i}^{\ast},S_{i},\theta_{0})]=0,\quad X_{i}=X_{i}^{\ast}+\varepsilon_{i},\quad\mathbb{E}[\varepsilon_{i}]=0.$ (G.1)
@@
Then, we can rewrite equation (G.1) as
@@
$\widetilde{\psi}(\widetilde{X}_{i},S_{i},\theta,\widetilde{\gamma})=\widetilde{g}(\widetilde{X}_{i},S_{i},\theta)-\sum_{k=2}^{K}\widetilde{\gamma}_{k}\widetilde{g}_{x}^{(k)}(\widetilde{X}_{i},S_{i},\theta),$ (G.2)
@@
Finally, let us now show that moment condition $\widetilde{\psi}$ in equation (G.2) is numerically identical to $\psi$ in equation (11) with $\gamma_{k}=\sigma_{X^{\ast}}^{k}\widetilde{\gamma}_{k}$. Note that for $\widetilde{x}=x/\sigma_{X^{\ast}}$ we have $\widetilde{g}_{\widetilde{x}}^{(k)}(\widetilde{x},s,\theta)\equiv\nabla_{\widetilde{x}}^{k}g(\sigma_{X^{\ast}}\widetilde{x},s,\theta)=\sigma_{X^{\ast}}^{k}g_{a}^{(k)}(a,s,\theta)|_{a=\sigma_{X^{\ast}}\widetilde{x}}=\sigma_{X^{\ast}}^{k}g_{x}^{(k)}(x,s,\theta)$, and hence
[…]
@@
Appendix H Some Implementation Details